OpenStudy (anonymous):

4log[12](2)+log[12](x)=log[12](96)

OpenStudy (anonymous):

start with \[\log_{12}(2^2)+\log_{12}(x)=\log_{12}(96)\] then go to \[\log_{12}(4x)=\log_{12}(96)\] and then it should be easy

OpenStudy (anonymous):

i think there's an error... it should be 2^4?

OpenStudy (anonymous):

\[\log_{12}(2^4)+\log_{12}(x)=\log_{12}(96)\]

OpenStudy (anonymous):

yes i made a mistake

OpenStudy (anonymous):

go with dpalnc answer

OpenStudy (anonymous):

\[\log_{12}(16x)=\log_{12}(96)\]

OpenStudy (anonymous):

from there can you solve for x?

OpenStudy (anonymous):

x=6 is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does 2log[2](2)+2log[2](6)-log[2](3x)=3 equal x=16??? thanks for the help